Find the probability that the number who have encountered fraudulent charges on their credit cards is fewer than 66. N=100, P=0.62
Answer #1

Answer From the provided information, Probability (P) = 0.62, sample size(o)=100 So, q=t-p = 1-0.62 = 0.38 The distributon wi